\begin{thebibliography}{35}
\providecommand{\natexlab}[1]{#1}
\providecommand{\url}[1]{\texttt{#1}}
\expandafter\ifx\csname urlstyle\endcsname\relax
  \providecommand{\doi}[1]{doi: #1}\else
  \providecommand{\doi}{doi: \begingroup \urlstyle{rm}\Url}\fi

\bibitem[Bannur et~al.(2023)Bannur, Hyland, Liu, Perez-Garcia, Ilse, Castro, Boecking, Sharma, Bouzid, Thieme, et~al.]{bannur2023learning}
Bannur, S., Hyland, S., Liu, Q., Perez-Garcia, F., Ilse, M., Castro, D.~C., Boecking, B., Sharma, H., Bouzid, K., Thieme, A., et~al.
\newblock Learning to exploit temporal structure for biomedical vision-language processing.
\newblock In \emph{Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition}, 2023.

\bibitem[Basu et~al.(2023)Basu, Massiceti, Hu, and Feizi]{basu2023strong}
Basu, S., Massiceti, D., Hu, S.~X., and Feizi, S.
\newblock Strong baselines for parameter efficient few-shot fine-tuning.
\newblock \emph{arXiv}, 2023.

\bibitem[Ben~Zaken et~al.(2022)Ben~Zaken, Goldberg, and Ravfogel]{bitfit}
Ben~Zaken, E., Goldberg, Y., and Ravfogel, S.
\newblock {B}it{F}it: Simple parameter-efficient fine-tuning for transformer-based masked language-models.
\newblock In \emph{ACL}, May 2022.

\bibitem[Butterick(2023)]{butterick_2023}
Butterick, M.
\newblock Stable diffusion litigation · joseph saveri law firm \& matthew butterick, Jan 2023.
\newblock URL \url{https://stablediffusionlitigation.com/}.

\bibitem[Carlini et~al.(2023)Carlini, Hayes, Nasr, Jagielski, Sehwag, Tramer, Balle, Ippolito, and Wallace]{carlini2023extracting}
Carlini, N., Hayes, J., Nasr, M., Jagielski, M., Sehwag, V., Tramer, F., Balle, B., Ippolito, D., and Wallace, E.
\newblock Extracting training data from diffusion models.
\newblock In \emph{32nd USENIX Security Symposium (USENIX Security 23)}, 2023.

\bibitem[Chambon et~al.(2022)Chambon, Bluethgen, Delbrouck, Van~der Sluijs, Po{\l}acin, Chaves, Abraham, Purohit, Langlotz, and Chaudhari]{chambon2022roentgen}
Chambon, P., Bluethgen, C., Delbrouck, J.-B., Van~der Sluijs, R., Po{\l}acin, M., Chaves, J. M.~Z., Abraham, T.~M., Purohit, S., Langlotz, C.~P., and Chaudhari, A.
\newblock Roentgen: Vision-language foundation model for chest x-ray generation.
\newblock \emph{arXiv preprint arXiv:2211.12737}, 2022.

\bibitem[Chen et~al.(2022)Chen, Ge, Tong, Wang, Song, Wang, and Luo]{chen2022adaptformer}
Chen, S., Ge, C., Tong, Z., Wang, J., Song, Y., Wang, J., and Luo, P.
\newblock Adaptformer: Adapting vision transformers for scalable visual recognition.
\newblock \emph{arXiv}, 2022.

\bibitem[Cohen et~al.(2020)Cohen, Hashir, Brooks, and Bertrand]{torchxrayvision}
Cohen, J.~P., Hashir, M., Brooks, R., and Bertrand, H.
\newblock On the limits of cross-domain generalization in automated x-ray prediction.
\newblock In \emph{Medical Imaging with Deep Learning}, 2020.
\newblock URL \url{https://arxiv.org/abs/2002.02497}.

\bibitem[Dhariwal \& Nichol(2021)Dhariwal and Nichol]{dhariwal2021diffusion}
Dhariwal, P. and Nichol, A.~Q.
\newblock Diffusion models beat {GAN}s on image synthesis.
\newblock In Beygelzimer, A., Dauphin, Y., Liang, P., and Vaughan, J.~W. (eds.), \emph{Advances in Neural Information Processing Systems}, 2021.
\newblock URL \url{https://openreview.net/forum?id=AAWuCvzaVt}.

\bibitem[Dutt et~al.(2023)Dutt, Ericsson, Sanchez, Tsaftaris, and Hospedales]{dutt2023parameter}
Dutt, R., Ericsson, L., Sanchez, P., Tsaftaris, S.~A., and Hospedales, T.
\newblock Parameter-efficient fine-tuning for medical image analysis: The missed opportunity.
\newblock \emph{arXiv}, 2023.

\bibitem[Dutt et~al.(2024)Dutt, Bohdal, Tsaftaris, and Hospedales]{dutt2024fairtune}
Dutt, R., Bohdal, O., Tsaftaris, S.~A., and Hospedales, T.
\newblock Fairtune: Optimizing parameter efficient fine tuning for fairness in medical image analysis.
\newblock In \emph{The Twelfth International Conference on Learning Representations}, 2024.
\newblock URL \url{https://openreview.net/forum?id=ArpwmicoYW}.

\bibitem[Han et~al.(2023)Han, Li, Zhang, Milanfar, Metaxas, and Yang]{han2023svdiff}
Han, L., Li, Y., Zhang, H., Milanfar, P., Metaxas, D., and Yang, F.
\newblock Svdiff: Compact parameter space for diffusion fine-tuning.
\newblock \emph{arXiv:2303.11305}, 2023.

\bibitem[Ho et~al.(2020)Ho, Jain, and Abbeel]{ho2020denoising}
Ho, J., Jain, A., and Abbeel, P.
\newblock Denoising diffusion probabilistic models.
\newblock In Larochelle, H., Ranzato, M., Hadsell, R., Balcan, M., and Lin, H. (eds.), \emph{Advances in Neural Information Processing Systems}, volume~33, pp.\  6840--6851. Curran Associates, Inc., 2020.
\newblock URL \url{https://proceedings.neurips.cc/paper_files/paper/2020/file/4c5bcfec8584af0d967f1ab10179ca4b-Paper.pdf}.

\bibitem[Hu et~al.(2022)Hu, yelong shen, Wallis, Allen-Zhu, Li, Wang, Wang, and Chen]{hu2022lora}
Hu, E.~J., yelong shen, Wallis, P., Allen-Zhu, Z., Li, Y., Wang, S., Wang, L., and Chen, W.
\newblock Lo{RA}: Low-rank adaptation of large language models.
\newblock In \emph{ICLR}, 2022.

\bibitem[Huang et~al.(2017)Huang, Liu, Van Der~Maaten, and Weinberger]{huang2017densely}
Huang, G., Liu, Z., Van Der~Maaten, L., and Weinberger, K.~Q.
\newblock Densely connected convolutional networks.
\newblock In \emph{Proceedings of the IEEE conference on computer vision and pattern recognition}, 2017.

\bibitem[Huang et~al.(2023)Huang, Huang, Yang, Ren, Liu, Li, Ye, Liu, Yin, and Zhao]{huang2023make}
Huang, R., Huang, J., Yang, D., Ren, Y., Liu, L., Li, M., Ye, Z., Liu, J., Yin, X., and Zhao, Z.
\newblock Make-an-audio: Text-to-audio generation with prompt-enhanced diffusion models.
\newblock In \emph{International Conference on Machine Learning}, pp.\  13916--13932. PMLR, 2023.

\bibitem[Johnson et~al.(2016)Johnson, Pollard, Shen, Lehman, Feng, Ghassemi, Moody, Szolovits, Anthony~Celi, and Mark]{johnson2016mimic}
Johnson, A.~E., Pollard, T.~J., Shen, L., Lehman, L.-w.~H., Feng, M., Ghassemi, M., Moody, B., Szolovits, P., Anthony~Celi, L., and Mark, R.~G.
\newblock Mimic-iii, a freely accessible critical care database.
\newblock \emph{Scientific data}, 2016.

\bibitem[Karras et~al.(2019)Karras, Laine, and Aila]{karras2019style}
Karras, T., Laine, S., and Aila, T.
\newblock A style-based generator architecture for generative adversarial networks.
\newblock In \emph{Proceedings of the IEEE/CVF conference on computer vision and pattern recognition}, 2019.

\bibitem[Kim et~al.(2022)Kim, Kim, and Yoon]{kim2022guided}
Kim, H., Kim, S., and Yoon, S.
\newblock Guided-tts: A diffusion model for text-to-speech via classifier guidance.
\newblock In \emph{International Conference on Machine Learning}, pp.\  11119--11133. PMLR, 2022.

\bibitem[Lian et~al.(2022)Lian, Zhou, Feng, and Wang]{lian2022scaling}
Lian, D., Zhou, D., Feng, J., and Wang, X.
\newblock Scaling \& shifting your features: A new baseline for efficient model tuning.
\newblock \emph{Advances in Neural Information Processing Systems}, 35:\penalty0 109--123, 2022.

\bibitem[Ramesh et~al.(2022)Ramesh, Dhariwal, Nichol, Chu, and Chen]{ramesh2022hierarchical}
Ramesh, A., Dhariwal, P., Nichol, A., Chu, C., and Chen, M.
\newblock Hierarchical text-conditional image generation with clip latents.
\newblock \emph{arXiv:2204.06125}, 2022.

\bibitem[Rombach et~al.(2022{\natexlab{a}})Rombach, Blattmann, Lorenz, Esser, and Ommer]{Rombach_2022_CVPR}
Rombach, R., Blattmann, A., Lorenz, D., Esser, P., and Ommer, B.
\newblock High-resolution image synthesis with latent diffusion models.
\newblock In \emph{Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR)}, pp.\  10684--10695, June 2022{\natexlab{a}}.

\bibitem[Rombach et~al.(2022{\natexlab{b}})Rombach, Blattmann, Lorenz, Esser, and Ommer]{rombach2022high}
Rombach, R., Blattmann, A., Lorenz, D., Esser, P., and Ommer, B.
\newblock High-resolution image synthesis with latent diffusion models.
\newblock In \emph{Proceedings of the IEEE/CVF conference on computer vision and pattern recognition}, pp.\  10684--10695, 2022{\natexlab{b}}.

\bibitem[Sohl-Dickstein et~al.(2015)Sohl-Dickstein, Weiss, Maheswaranathan, and Ganguli]{sohl2015deep}
Sohl-Dickstein, J., Weiss, E., Maheswaranathan, N., and Ganguli, S.
\newblock Deep unsupervised learning using nonequilibrium thermodynamics.
\newblock In Bach, F. and Blei, D. (eds.), \emph{Proceedings of the 32nd International Conference on Machine Learning}, volume~37 of \emph{Proceedings of Machine Learning Research}, pp.\  2256--2265, Lille, France, 07--09 Jul 2015. PMLR.

\bibitem[Somepalli et~al.(2023{\natexlab{a}})Somepalli, Singla, Goldblum, Geiping, and Goldstein]{somepalli2023diffusion}
Somepalli, G., Singla, V., Goldblum, M., Geiping, J., and Goldstein, T.
\newblock Diffusion art or digital forgery? investigating data replication in diffusion models.
\newblock In \emph{Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition}, 2023{\natexlab{a}}.

\bibitem[Somepalli et~al.(2023{\natexlab{b}})Somepalli, Singla, Goldblum, Geiping, and Goldstein]{somepalli2023understanding}
Somepalli, G., Singla, V., Goldblum, M., Geiping, J., and Goldstein, T.
\newblock Understanding and mitigating copying in diffusion models.
\newblock \emph{Advances in Neural Information Processing Systems}, 36, 2023{\natexlab{b}}.

\bibitem[Song et~al.(2021{\natexlab{a}})Song, Meng, and Ermon]{song2021denoising}
Song, J., Meng, C., and Ermon, S.
\newblock Denoising diffusion implicit models.
\newblock In \emph{International Conference on Learning Representations}, 2021{\natexlab{a}}.
\newblock URL \url{https://openreview.net/forum?id=St1giarCHLP}.

\bibitem[Song et~al.(2021{\natexlab{b}})Song, Sohl-Dickstein, Kingma, Kumar, Ermon, and Poole]{song2021scorebased}
Song, Y., Sohl-Dickstein, J., Kingma, D.~P., Kumar, A., Ermon, S., and Poole, B.
\newblock Score-based generative modeling through stochastic differential equations.
\newblock In \emph{International Conference on Learning Representations}, 2021{\natexlab{b}}.
\newblock URL \url{https://openreview.net/forum?id=PxTIG12RRHS}.

\bibitem[Touvron et~al.(2022)Touvron, Cord, El-Nouby, Verbeek, and J{\'e}gou]{touvron2022three}
Touvron, H., Cord, M., El-Nouby, A., Verbeek, J., and J{\'e}gou, H.
\newblock Three things everyone should know about vision transformers.
\newblock In \emph{ECCV}, 2022.

\bibitem[Vignac et~al.(2023)Vignac, Krawczuk, Siraudin, Wang, Cevher, and Frossard]{vignac2023digress}
Vignac, C., Krawczuk, I., Siraudin, A., Wang, B., Cevher, V., and Frossard, P.
\newblock Digress: Discrete denoising diffusion for graph generation.
\newblock In \emph{The Eleventh International Conference on Learning Representations}, 2023.
\newblock URL \url{https://openreview.net/forum?id=UaAD-Nu86WX}.

\bibitem[von Platen et~al.(2022)von Platen, Patil, Lozhkov, Cuenca, Lambert, Rasul, Davaadorj, Nair, Paul, Berman, Xu, Liu, and Wolf]{von-platen-etal-2022-diffusers}
von Platen, P., Patil, S., Lozhkov, A., Cuenca, P., Lambert, N., Rasul, K., Davaadorj, M., Nair, D., Paul, S., Berman, W., Xu, Y., Liu, S., and Wolf, T.
\newblock Diffusers: State-of-the-art diffusion models.
\newblock \url{https://github.com/huggingface/diffusers}, 2022.

\bibitem[Webster et~al.(2023)Webster, Rabin, Simon, and Jurie]{webster2023duplication}
Webster, R., Rabin, J., Simon, L., and Jurie, F.
\newblock On the de-duplication of laion-2b.
\newblock \emph{arXiv:2303.12733}, 2023.

\bibitem[Wen et~al.(2024)Wen, Liu, Chen, and Lyu]{wen2024detecting}
Wen, Y., Liu, Y., Chen, C., and Lyu, L.
\newblock Detecting, explaining, and mitigating memorization in diffusion models.
\newblock In \emph{The Twelfth International Conference on Learning Representations}, 2024.
\newblock URL \url{https://openreview.net/forum?id=84n3UwkH7b}.

\bibitem[Xie et~al.(2023)Xie, Yao, Shi, Liu, Zhou, Liu, Li, and Li]{xie2023difffit}
Xie, E., Yao, L., Shi, H., Liu, Z., Zhou, D., Liu, Z., Li, J., and Li, Z.
\newblock Difffit: Unlocking transferability of large diffusion models via simple parameter-efficient fine-tuning.
\newblock \emph{arXiv:2304.06648}, 2023.

\bibitem[Xu et~al.(2022)Xu, Yu, Song, Shi, Ermon, and Tang]{xu2022geodiff}
Xu, M., Yu, L., Song, Y., Shi, C., Ermon, S., and Tang, J.
\newblock Geodiff: A geometric diffusion model for molecular conformation generation.
\newblock In \emph{International Conference on Learning Representations}, 2022.
\newblock URL \url{https://openreview.net/forum?id=PzcvxEMzvQC}.

\end{thebibliography}

\documentclass{article}

\usepackage{microtype}
\usepackage{graphicx}
\usepackage{subfigure}
\usepackage{booktabs} 

\usepackage{hyperref}
\usepackage{colortbl}
\usepackage{amsmath}

\usepackage[accepted]{icml2024}

\usepackage{amsmath}
\usepackage{amssymb}
\usepackage{mathtools}
\usepackage{amsthm}

\usepackage[capitalize,noabbrev]{cleveref}

\theoremstyle{plain}

\theoremstyle{definition}

\theoremstyle{remark}

\usepackage[textsize=tiny]{todonotes}

\icmltitlerunning{Capacity Control is an Effective Memorization Mitigation Mechanism}

\DeclareMathOperator{\argmin}{arg\,min}

\begin{document}

\twocolumn[
\icmltitle{Capacity Control is an Effective Memorization Mitigation Mechanism in Text-Conditional Diffusion Models}

\icmlsetsymbol{equal}{*}

\begin{icmlauthorlist}
\icmlauthor{Raman Dutt}{yyy}
\icmlauthor{Pedro Sanchez}{xxx}
\icmlauthor{Ondrej Bohdal}{yyy}
\icmlauthor{Sotirios A. Tsaftaris }{xxx}
\icmlauthor{Timothy Hospedales}{yyy,comp}
\end{icmlauthorlist}

\icmlaffiliation{yyy}{School of Informatics, University of Edinburgh, Edinburgh, UK}
\icmlaffiliation{xxx}{School of Engineering, University of Edinburgh, Edinburgh, UK}
\icmlaffiliation{comp}{Samsung AI Center, Cambridge, UK}

\icmlcorrespondingauthor{Raman Dutt}{raman.dutt@ed.ac.uk}

\icmlkeywords{Diffusion Models, Memorization, PEFT, Medical Image Analysis}

\vskip 0.3in
]

\printAffiliationsAndNotice{\icmlEqualContribution} 



\section{Introduction and Related Work}
\label{submission}

Generative diffusion models \cite{sohl2015deep, song2021denoising, ho2020denoising, song2021scorebased} excel in generating high-quality data across various formats, including images \cite{dhariwal2021diffusion, karras2019style, ramesh2022hierarchical, rombach2022high}, audio \cite{kim2022guided, huang2023make}, and graphs \cite{xu2022geodiff, vignac2023digress}. These models power successful applications like \textit{Midjourney}, \textit{Stable Diffusion}, and \textit{DALL-E}. The strength of these models lies in their ability to produce high-quality, novel images that capture real-world complexities. A notable application is using synthetic data to share sensitive information without privacy risks, thus facilitating progress. However, this benefit is compromised if models inadvertently reproduce training data, known as \textit{memorization}, raising legal and ethical concerns \cite{butterick_2023}. This issue has prompted the development of mitigation methods, including training-time interventions  \cite{webster2023duplication, somepalli2023understanding} noise regularizers \cite{somepalli2023diffusion}, and inference-time solutions \cite{somepalli2023understanding, wen2024detecting}.

In domain-specific diffusion models for medical imaging, the process typically involves starting with a general-purpose model and fine-tuning it with specific data \cite{chambon2022roentgen}. Our paper hypothesizes that the overcapacity of large neural networks leads to memorization. We explore regularizing model capacity during fine-tuning to mitigate this issue, using Parameter-Efficient Fine-Tuning (PEFT) methods from text-conditional diffusion models. Our results show that capacity-controlled approaches, like PEFT, effectively mitigate memorization during training in diffusion models.

\section{Capacity Control using Parameter-Efficient Fine-Tuning}

The conventional way of adapting a \emph{pre-trained} model to a specific domain is by updating the entire parameter set using a downstream dataset $\mathcal{D}^{train}$ and a task-specific loss $\mathcal{L}^{base}$. An alternate is offered by Parameter-Efficient Fine-Tuning (PEFT) methods that involve freezing most of the original \emph{pre-trained} model and fine-tuning a specific parameter subset. PEFT methods can be broadly categorized into \textit{Selective} and \textit{Additive} strategies. The \textbf{Selective} category includes methods that involve fine-tuning a subset of existing model parameters such as attention matrices \cite{touvron2022three}, normalization layers \cite{basu2023strong} and bias parameters \cite{bitfit}. The \textbf{Additive} category of methods freezes the original model and introduces a new set of parameters for fine-tuning \cite{lian2022scaling,chen2022adaptformer,han2023svdiff,hu2022lora,xie2023difffit}. For text-to-image generation, strategies such as SV-DIFF \cite{han2023svdiff} and DiffFit \cite{xie2023difffit} have been proposed, and they excel in both in-domain and out-of-domain settings \cite{dutt2023parameter}. 

The key idea of PEFT is to fine-tune a very small subset of model parameters $\phi \subset \theta$, i.e. $|\phi| \ll |\theta|$. The parameters of the pre-trained model are initially $\theta_0$, and after fine-tuning on data $\mathcal{D}^{train}$ they change by $\Delta\phi$ compared to their initial values. The fine-tuning process can then be defined via:


\begin{equation}
    \Delta\phi^* = \argmin_{\Delta\phi} \mathcal{L}^{base} \left( \mathcal{D}^{train}; \theta_0 + \omega \odot \Delta\phi  \right)
\end{equation}

where $\mathcal{L}^{base}$ is the mean squared error in the context of image generation problems.

\section{Experiments}

\subsection{Experimental Setup}
\textbf{Architecture:} Our experiments adopted the pre-trained Stable Diffusion (v1.5) \cite{Rombach_2022_CVPR} model implemented in the \textit{Diffusers} package \cite{von-platen-etal-2022-diffusers}. For fine-tuning on medical data we follow previous practice \cite{dutt2023parameter, chambon2022roentgen}, and fine-tune only the U-Net while keeping the other components in the stable diffusion pipeline frozen during our experiments. 

\textbf{Datasets: } We use the MIMIC dataset \cite{johnson2016mimic} that consists of chest X-rays and associated radiology text reports. The dataset was split into train, validation and test splits ensuring no leakage of subjects, resulting in 110K training samples, 14K validation and 14K test samples. Note that several images in this dataset can be associated with the same text prompt due to similarity in clinical findings.


\subsection{Memorization and Image-Quality Metrics} \label{sec:metrics}

\textbf{Memorization: Nearest Neighbour Distance}\quad A generated example $\hat{x}$ is considered memorized if it is very close to a training example $x$, measured by a distance metric $l$ with threshold $\delta$. We report the average minimum cosine distance between each synthetic sample and its nearest training example.
\textbf{Memorization: Extraction Attack}\quad We assess memorization using an image extraction attack \cite{carlini2023extracting}, which involves generating many images for a given prompt and performing membership inference to identify memorized images. Fewer extracted images indicate less memorization.
\textbf{Memorization: Denoising Strength Heuristic}\quad As a proxy for memorization, we use the magnitude of the diffusion noise prediction \cite{wen2024detecting}. This measures the difference between model predictions conditioned on a prompt embedding $e_{p}$ and an empty string embedding $e_{\phi}$ averaged over multiple time steps.
\textbf{Quality: Fréchet Inception Distance (FID)}\quad FID score evaluates image generation quality by comparing distributions of real and synthetic images based on features from a pre-trained image encoder. We use a DenseNet-121 model \cite{huang2017densely} pre-trained on chest X-rays \cite{torchxrayvision} for feature extraction.
\textbf{Quality: BioViL-T Score}\quad BioViL-T \cite{bannur2023learning}, a vision-language model for chest X-rays and radiology reports, provides a joint score indicating the clinical correlation between an image and a text caption. Higher scores reflect better quality and textual guidance of generated images.

\section{Results}

In this section, we report our results from the experiments performed on the MIMIC dataset for generation quality (FID $\downarrow$, BioViL-T Score $\uparrow$) and memorization (AMD $\uparrow$, Extracted Images $\downarrow$).

\textbf{Table \ref{tab:all_results_combined}} reports the results for comparing PEFT strategies with conventional fine-tuning methods. PEFT methods act as strong regularizers and hence, significantly restrict the model capacity during fine-tuning. The results indicate that controlling the model capacity can enable both high-generation quality and memorization mitigation. 

\textbf{Table \ref{tab:table2}} indicates results for combining different fine-tuning methods (conventional and PEFT) with existing memorization mitigation strategies \cite{somepalli2023understanding,wen2024detecting}. The results indicate that while existing mitigation strategies do play their part, they can easily be combined with PEFT methods to reduce memorization further and significantly outperform conventional fine-tuning.

\begin{table}[ht]
\centering
\caption{Comparitive evaluation of generation quality (FID $\downarrow$, BioViL-T Score $\uparrow$) and memorization (AMD $\uparrow$, Extracted Images $\downarrow$) on the MIMIC dataset. PEFT strategies consistently outperform conventional fine-tuning methods in terms of both generation quality and memorization.}
\resizebox{0.5\textwidth}{!}{%
\begin{tabular}{@{}lcccc@{}}
\toprule
\textbf{FT Method}                  & \textbf{FID Score ($\downarrow$)} & \textbf{AMD ($\uparrow$)} & \textbf{BioViL-T Score ($\uparrow$)} & \textbf{Extracted Images ($\downarrow$)} \\ \midrule
Freeze                             & 323.78                            & \textbf{0.801}                        & 0.23             & \phantom{00}0                    \\
Full FT                            & \phantom{0}35.79                             & 0.029                       & 0.60             & 356                    \\


DiffFit                          & \phantom{0}15.19                             & 0.061                        & 0.69       & \phantom{0}69                          \\
Attn Tuning          & \phantom{0}40.28                             & 0.031                       & 0.59       & 358                          \\
SV-DIFF   & \phantom{0}\textbf{12.77}                             & 0.062                        & \textbf{0.72}       & \phantom{0}\textbf{63}                        \\

 \bottomrule
\end{tabular}
}
\label{tab:all_results_combined}
\end{table}

\begin{table}[ht]
\centering
\caption{Combining different fine-tuning strategies combined with existing memorization mitigation mechanisms.}
\resizebox{0.5\textwidth}{!}{%
\begin{tabular}{@{}lcccc@{}}
\toprule
\textbf{FT Method}                  & \textbf{FID Score ($\downarrow$)} & \textbf{AMD ($\uparrow$)} & \textbf{BioViL-T Score ($\uparrow$)} & \textbf{Extracted Images ($\downarrow$)} \\ \midrule

Full FT  + RWA                      & \phantom{0}67.48                             & 0.041                       & 0.58       & 320                          \\
Full FT  + Threshold Mitigation     & \phantom{0}25.43                             & 0.048                        & 0.62       & 298                           \\ \\

DiffFit  RWA                     & \phantom{0}17.18                             & 0.060                        & 0.67       & \phantom{0}64                          \\
DiffFit  Threshold Mitigation     & \phantom{0}\textbf{15.31}                             & 0.067                        & \textbf{0.69}       & \phantom{0}51                          \\ \\

Attn Tuning  + RWA                 & \phantom{0}49.71                             & 0.035                       & 0.52       & 324                          \\
Attn Tuning  + Threshold Mitigation   & \phantom{0}25.43                             & 0.041                        & 0.62       & 299                          \\ \\
SV-DIFF  + RWA                     & \phantom{0}41.25                             & 0.059                        & 0.52       & \phantom{0}58                          \\
SV-DIFF  + Threshold Mitigation      & \phantom{0}23.64                             & \textbf{0.068}                       & 0.64       & \phantom{0}\textbf{45}                          \\

 \bottomrule
\end{tabular}
}
\label{tab:table2}
\end{table}



\section{Conclusion and Future Work}

In this work, we addressed memorization in diffusion models by controlling the model's capacity. Our experiments showed that model capacity significantly influences memorization. By employing parameter-efficient fine-tuning (PEFT), we were able to reduce memorization while maintaining high-generation quality. Additionally, we demonstrated that these fine-tuning methods can complement existing memorization mitigation strategies.

For future work, we aim to develop an automated scheme for selecting the optimal PEFT parameters during fine-tuning to balance memorization and generalization. Similar frameworks have been proposed for optimizing parameter updates based on reward metrics \cite{dutt2024fairtune}. We plan to formulate our approach as a bi-level optimization problem, where the outer loop proposes parameter subsets to update, and the inner loop fine-tunes these parameters on the downstream task. The process will be guided by reward metrics that assess both generation quality and memorization. In this process, our framework can find the optimal model capacity (parameter subsets) to update to improve both generation quality and reduce memorization.

\bibliography{example_paper}
\bibliographystyle{icml2024}

\newpage
\appendix
\onecolumn

\end{document}